A random sample of 9 history students produced the following data,

first test score, x second test score, y
90 51
36 82
13 15
71 77
82 10
20 18
92 35
53 49
63 85

where x measures the first test score and y measures the second test score. What is the estimate of the y = a x + b regression slope coefficient?
Answer #1

Here,

X Y
90 51
36 82
13 15
71 77
82 10
20 18
92 35
53 49
63 85

From the data, we have,

n=9,\sum xy=25556,\overline{x}=57.7778,\overline{y}=46.8889,\sum x^2=36972

The regression slope is given by:

a=\frac{\sum xy-n\overline{x}.\overline{y}}{\sum x^2-n\overline{x}^2} =\frac{25556-9(57.7778)(46.8889)}{36792-9(57.7778)^2}=0.17395